# Discrete Heat Kernels on Simplicial Complexes and Its Application to Functional Brain Networks

### v1

@@ 1 Introduction
Networks embeds the interactions of elements in a physical or virtual space. Graphs are often used to represent and describe the geometric structures of network data, and in their basic form are composed of nodes or data points in which for any two pair of nodes, there exists a possible edge between them governed by a criteria, such as correlation. The nodes and edges in a graph are often endowed with a weighting function that defines edge weights. These weights depict signals on the graph. Modeling signals on graphs falls under the common research paradigm of graph signal processing [36, 39], in which classical signal processing methodologies are extended to irregular data structures such as Graphs.
@@
Graph signals are ubiquitous in many biological phenomena. In brain imaging, non-invasive procedures have been devised to segregate functional regions of the human cerebral cortex which are interconnected by a dense network of cortico-cortical axonal pathways [27, 33]. Anatomical surfaces, extracted from magnetic resonance images (MRI) and computed tomography (CT) are usually represented as triangle meshes, a special case of graphs [12]. Such mesh data are usually noisy. In such graph data, measurements such as cortical thickness and surface curvature are obtained at mesh vertices, which form the nodes of the graph. In neuroscience, the interconnectivity of the axonal pathways can in themselves be represented by a graph object [9].
@@
Often signal processing techniques are applied to such data. Processing graph signals takes various forms; smoothing, denoising, filtering among others. It has had quite successful applications in neuroimaging. Diffusion wavelets have been applied to characterize the localized growth patterns of mandible surfaces obtained in CT images, while graph signal processing can be applied to filter brain activity based on concepts of spectral modes derived from brain structure, and can be extended to brain imaging classification task [29, 31].
@@
Despite this success, signal processing on graph structures have the limitation that, it captures only pairwise interactions. Many physical and biological systems encodes more complex structures beyond pairwise interactions. Few studies in neuroimaging and biological systems have explored higher-order interactions in networks [14, 15, 22, 38, 52]. The majority of methods try to extend and generalize the pairwise relationships to group-level interactions. However, this generalization in itself is limiting, as there are not explicit in the representation of the true underlying structure of higher dimensional networks.
@@
Fig. 1: A schematic of signals on a simplicial complex of different orders. (A) Signals on 0-simplices. (B) Signals on 1-simplices and (C) Signals on 2-simplices $\{[1,2],[2,3],[1,2]\},\{[1,3],[1,4],[3,5]\}$. Arrows depicts the chosen orientation.
@@
Simplicial complexes are algebraic objects often studied in persistent homology (PH) and topological data analysis (TDA) as the basic building block for representing complex data (see Section 2 for more details). Simplicial complex representation has been extensively applied to brain network and neuroimaging task. The encoding of the structural and functional connections of brain neurons using simplicial complexes has been studied [3, 9, 14], and methodologies developed for exploring their statistical properties in this representation [17, 25]. Signal processing on higher order simplices, unlike their lower order counterparts like graphs is not that straightforward. Recognizing the notion of flow in edges (Figure 1), which is non-existent in signal processing over vertices, the concept of Edge-Laplacian, a special case of the Hodge-Laplacian was devised to study signals over edges [45]. A general framework for signal processing over complexes was established, with methods developed to handle only the case of signals defined over edges , with a similar exposition on simplicial complexes as a modeling framework for higher-order network interactions, signal smoothing and denoising via Fourier transform [5, 45]. Detail methods and analysis, however, were limited to the special case of $1$-simplices.
@@
A popular algebraic tool for representing topological features in high dimensions is the persistent homology (PH)[18]. This allows for the computation of topological features of a space (represented as a simplicial complex) at different spatial resolutions. Usually this computation is preceded by imposing a function on the underlying space, which acts on the space to generate a nested sequence of simplicial complexes. The resulting nested sequence of high-order networks is referred to as a filtration. In medical imaging and other biological applications, PH has been applied to establish algebraic and statistical properties of various topological features in brain networks, establish statistical methods for studying electroencephalographic signals, analyzing breathing sound signals, and arterial pulse pressure waves among others [19, 20]. The central theme in all these works is to develop inference procedures on topological features that persists over wide spatial scales, which are likely not to represent noise. One topological invariant that is used to distinguish topological spaces is connectedness. This invariant represents the integration of brain networks. Cycles, a second topological invariant measures the strength of this integration. Whiles brain connectivity modeled as 1-skeleton has been extensively studied in the literature[11, 40], cycles has not enjoyed the same exploration mainly due to the fact that it’s difficult to compute even for the simple case of 1-skeletons [11, 15]. Computing cycles in 1-skeletons is often done using brute force approach [48]. More recently, a scalable algorithm for computing cycles in 1-skeletons and a method for computing and identifying cycles in high-order topological spaces via the Hodge Laplacian has been developed [2, 14]. This provides an efficient algorithm for extracting cycles that retains the simplicity of graph-based approaches.
@@ 2.1 Graphs
A graph is an ordered set $\mathcal{X}=\left(V(\mathcal{X}),E(\mathcal{X}),\psi_{\mathcal{X}}\right)$, where $V(\mathcal{X})$ is the set of nodes, $E(\mathcal{X})$ is the set of edges and $\psi_{\mathcal{X}}$ is an incidence function connecting any pair of nodes of $V(\mathcal{X})$ to form edges $E(\mathcal{X})$. When the underlying graph is unambiguous from the context we usually denote the node set and vertex set as $V$ and $E$ respectively. The number of nodes and edges are also denoted as $v(\mathcal{X})$ (also denoted as p) and $e(\mathcal{X})$ respectively. Two nodes incident by the same edge are said to be adjacent, similarly two edges incident by the same node are also adjacent [8]. They are two important matrices associated with graphs; the adjacency matrix $\textbf{A}(\mathcal{X})$ which is a $v(\mathcal{X})\times v(\mathcal{X})$} matrix with the $a_{ij}]$ entry been the number of edges connecting node $v_{i}$ and $v_{j}$, and the incidence matrix $\textbf{B}(\mathcal{X})$ which is a $v(\mathcal{X})\times e(\mathcal{X})$ matrix with the $b_{ij}]$ entry been the number of times node $v_{i}$ is incident with edge $e_{j}$, with loops contributing two (2) [8, 26]. $\textbf{B}(\mathcal{X})$ is an alternate form of depicting a graph and we can derive the degree of a node by taking the appropriate column sum of $\textbf{A}(\mathcal{X})$.
@@ 2.2 Simplicial Complex and Simplicial k-chains:
A k-simplex is a k-dimensional polytope whose convex hull is made up of its k+1 nodes. A simplicial complex $\mathcal{K}$ is a set of simplices such that for any $\sigma_{1},\sigma_{2}\in\mathcal{K}$, $\sigma_{1}\cap\sigma_{2}$ is a face of both simplices; and a face of any simplex $\sigma\in\mathcal{K}$ is also a simplex in $\mathcal{K}$.
For some $r_{i}\in\mathbb{R}$, the the formal sum $\sum_{i=1}^{N}r_{i}\sigma_{i}$, $\sigma_{i}\in\mathcal{K}$ is termed the simplicial k-chain. The set of simplicial k-chains with formal addition over $\mathbb{R}$ is known as an $\mathbb{R}$-module, and is used to construct boundary maps [49].
@@ 2.2.1 Boundary Map of Simplicial k-chains:
where $(v_{0},...,\hat{v_{i}},...,v_{k})$ is the k-1 face of $\sigma$ obtained by deleting the $v_{i}$ node.
The boundary map satisfies the property that $\partial_{k}o\partial_{k+1}=0$.
@@ 2.2.2 Homology of a Simplicial Complex $\mathcal{K}$:
Two important components of the boundary map (2) are it’s kernel denoted $\mathcal{Z}_{k}=ker(\partial_{k})$, and it’s image denoted $\mathcal{B}_{k}=img(\partial_{k+1})$. These are subspaces of $\mathcal{K}_{k}$. The elements of $\mathcal{Z}_{k}$ are known as $k$-cycles and that of $\mathcal{B}_{k}$ are known as $k$-boundaries [28]. In essence $\mathcal{Z}_{k}$ is the subspace of $\mathcal{K}_{k}$ consisting of $k$-chains that are also $k$-cycles and $\mathcal{B}_{k}$ is a subspace of $\mathcal{K}_{k}$ consisting of $k$-cycles that are also $k$-boundaries.
The set quotient $\mathcal{H}_{k}=\mathcal{Z}_{k}/\mathcal{B}_{k}$ is termed the kthkth homology-module and its an $\mathbb{R}$-module. It can be shown that $\mathcal{B}_{k}\subseteq\mathcal{Z}_{k}$ [28, 49]. Hence the difference, can intuitively be thought of as the failure of k-cycles in $\mathcal{K}$ to bound $(k+1)$-simplices, leading to the concept of ”holes”. A measure of this number of $k$-dimensional holes is termed the Betti number denoted $\beta_{k}=rank(\mathcal{H}_{k})$.
@@ 2.3 Homology of Graphs:
Graphs have natural relationship with complexes. An unoriented graph can be viewed as a 1-dimensional simplicial complex, and higher dimensional simplicial complexes are referred to as hypergraphs [32].
Consider the following characterization of a graph $\mathcal{X}=(V,E,w)$. Define $\mathcal{V}$ and $\mathcal{E}$ to be the node space and edge space of the graph. The formal sum over the node space
@@
Since $img(\partial_{2})=0$ [47], we have that the first homology module $\mathcal{H}_{1}\mathcal{(X)}=\ker(\partial_{1})$. If $\mathcal{X}$ is finite, then $\mathcal{H}_{1}\mathcal{(X)}$ has finite rank, given by the first Betti number $\beta_{1}$. Elements of $\mathcal{H}_{1}\mathcal{(X)}$ are termed 1-cycles.
@@
Let the complete graph $\mathcal{X}$ has the form $\mathcal{X}=(V,E,w)$. Since $\mathcal{X}$ is complete, it’s also connected with 1 connected component. The $0th$ Betti number $\beta_{0}$ is the number of connected components of $\mathcal{X}$. From the discussion in homology of graphs, the number of 1-cycles will be given by the $1st$ Betti number $\beta_{1}$. The Euler characteristic function of $\mathcal{X}$ is given by $\chi(\mathcal{X})=|V|-|E|=\beta_{0}-\beta_{1}$ [1].
The graph $\mathcal{X}$ is complete $\implies$ $|V|=p$, $|E|=\frac{|V|(|V|-1)}{2}=\frac{p(p-1)}{2}$. Hence
@@ 2.3.1 Graph Filtration
Intuitively, since $\mathcal{X}_{0}$ is complete, the number of connected components is therefore 1, as we increase threshold $\epsilon$, more edges are disconnected, increasing the number of connected components $(\beta_{0})$, and decreasing the number of cycles $(\beta_{1})$. More formally, it has been proved that in a graph $\mathcal{X}$, Betti numbers $\beta_{0}$ and $\beta_{1}$ are monotone over graph filtration on edge weights [46].
@@
[1]. Note that $\beta_{0}(\mathcal{X}_{\epsilon})=_{0}(\mathcal{X}_{1{(w_{ij}>\epsilon)}})$ and $|V|$ is the number of nodes in $\mathcal{X}_{\epsilon}$. The function $f(|V|,\epsilon)$ is a decreasing count of the edge set over the filtration. Since the first Betti number measures the number of 1-cycles, rearranging we have:
@@ 2.3.2 Betti Curves of Graph Filtration
Given the defined weights on $\mathcal{X}$, we can induce a filtration by appealing to a graph descriptor function. A graph descriptor function function $g$ is a mapping from $\mathcal{X}$ to some $d$-dimensional space, $g:\mathcal{X}\to\mathbb{R}^{d}$. That’s, it evaluates certain attributes of $\mathcal{X}$ and embeds them in some $d$-dimensional space.
For each of the subgraphs in (4), we can construct a higher dimensional path
@@
Fig. 2: A schematic of a sample filtration sequence of a seven node network $\mathcal{X}$. A weighted graph (far left) is thresholded at an increasing sequence of filtration values. A graph descriptor function $g$ (count of the number of connected components) is evaluated at each filtration value. The thresholding results in a sequence of nested graphs as shown in (4). The collection of such nested sequence of graphs defines a filtration.
@@ 2.3.3 Random Graphs
Random graphs construction are mostly inspired by the Erdos-Renyi (ER) random graph [7]. They are two theoretical approaches to generating random graphs based of the ER model. The first approach generates a random graph with a specified set of edges, say $M$, with $v$ nodes by randomly choosing an $M$-edge graph from the subset $\binom{N}{M}$ $\frac{N!}{(N-M)!M!}$, where $N=\frac{1}{2}p(p-1)$ is the maximum number of edges in a $p$-node graph. The second approach, which is more commonly used involves assigning an edge placement probability $\pi$ to each of the $N$ possible edges. For example, assume we want to generate a random graph of order $p$ (number of nodes), we will choose each of the $N$ potential edges, run a random number generator (or equivalently flip a coin), that outputs an integer say 1 with probability $\pi$, indicating that the edge should be created, or not created otherwise. The probability space associated with this second generation process has lot of desirable properties and has been extensively studied [21, 24].
Graphs modelled after ER are often unweighted. However, complex networks will require the characterization of the heterogeneity present in edge weights. The construction of weighted random graphs has been studied as an extension of the ER model [23], as a generalization of the unweighted graph and also by looking at an ensemble of weighted graphs [4]. In it’s simplest form, once a random edge weight has been placed, a random weight (often modelled after a probability distribution) is assigned to the edge.

### v2

@@ 1 Introduction
Networks embeds the interactions of elements in a physical or virtual space. Graphs are often used to represent and describe the geometric structures of network data, and in their basic form are composed of nodes or data points in which for any two pair of nodes, there exists a possible edge between them governed by a criteria, such as correlation. The nodes and edges in a graph are often endowed with a weighting function that defines edge weights. These weights depict signals on the graph. Modeling signals on graphs falls under the common research paradigm of graph signal processing [newman2018networks, ortega2018graph], in which classical signal processing methodologies are extended to irregular data structures such as Graphs.
@@
Graph signals are ubiquitous in many biological phenomena. In brain imaging, non-invasive procedures have been devised to segregate functional regions of the human cerebral cortex which are interconnected by a dense network of cortico-cortical axonal pathways [hagmann2008mapping, macdonald2000automated]. Anatomical surfaces, extracted from magnetic resonance images (MRI) and computed tomography (CT) are usually represented as triangle meshes, a special case of graphs [chung2015unified]. Such mesh data are usually noisy. In such graph data, measurements such as cortical thickness and surface curvature are obtained at mesh vertices, which form the nodes of the graph. In neuroscience, the interconnectivity of the axonal pathways can in themselves be represented by a graph object [bullmore2009complex].
@@
Often signal processing techniques are applied to such data. Processing graph signals takes various forms; smoothing, denoising, filtering among others. It has had quite successful applications in neuroimaging. Diffusion wavelets have been applied to characterize the localized growth patterns of mandible surfaces obtained in CT images, while graph signal processing can be applied to filter brain activity based on concepts of spectral modes derived from brain structure, and can be extended to brain imaging classification task [hu2016matched, huang2018graph].
@@
Despite this success, signal processing on graph structures have the limitation that, it captures only pairwise interactions. Many physical and biological systems encodes more complex structures beyond pairwise interactions. Few studies in neuroimaging and biological systems have explored higher-order interactions in networks [dakurah2022modelling, dakurahregistration, ganmor2011sparse, ohiorhenuan2010sparse, yu2011higher]. The majority of methods try to extend and generalize the pairwise relationships to group-level interactions. However, this generalization in itself is limiting, as there are not explicit in the representation of the true underlying structure of higher dimensional networks.
@@
Figure 1: A schematic of signals on a simplicial complex of different orders. (A) Signals on 0-simplices. (B) Signals on 1-simplices and (C) Signals on 2-simplices $\{[1,2],[2,3],[1,2]\},\{[1,3],[1,4],[3,5]\}$. Arrows depicts the chosen orientation.
@@
Simplicial complexes are algebraic objects often studied in persistent homology (PH) and topological data analysis (TDA) as the basic building block for representing complex data (see Section 2 for more details). Simplicial complex representation has been extensively applied to brain network and neuroimaging task. The encoding of the structural and functional connections of brain neurons using simplicial complexes has been studied [andjelkovic2020topology, bullmore2009complex, dakurah2022modelling], and methodologies developed for exploring their statistical properties in this representation [dakurah2025maxtda, giusti2016two]. Signal processing on higher order simplices, unlike their lower order counterparts like graphs is not that straightforward. Recognizing the notion of flow in edges (Figure 1), which is non-existent in signal processing over vertices, the concept of Edge-Laplacian, a special case of the Hodge-Laplacian was devised to study signals over edges [schaub2018flow]. A general framework for signal processing over complexes was established, with methods developed to handle only the case of signals defined over edges , with a similar exposition on simplicial complexes as a modeling framework for higher-order network interactions, signal smoothing and denoising via Fourier transform [barbarossa2020topological, schaub2018flow]. Detail methods and analysis, however, were limited to the special case of $1$-simplices.
@@
A popular algebraic tool for representing topological features in high dimensions is the persistent homology (PH)[edelsbrunner2008persistent]. This allows for the computation of topological features of a space (represented as a simplicial complex) at different spatial resolutions. Usually this computation is preceded by imposing a function on the underlying space, which acts on the space to generate a nested sequence of simplicial complexes. The resulting nested sequence of high-order networks is referred to as a filtration. In medical imaging and other biological applications, PH has been applied to establish algebraic and statistical properties of various topological features in brain networks, establish statistical methods for studying electroencephalographic signals, analyzing breathing sound signals, and arterial pulse pressure waves among others [emrani2014persistent, emrani2015novel]. The central theme in all these works is to develop inference procedures on topological features that persists over wide spatial scales, which are likely not to represent noise. One topological invariant that is used to distinguish topological spaces is connectedness. This invariant represents the integration of brain networks. Cycles, a second topological invariant measures the strength of this integration. Whiles brain connectivity modeled as 1-skeleton has been extensively studied in the literature[chung2019statistical, park2013structural], cycles has not enjoyed the same exploration mainly due to the fact that it’s difficult to compute even for the simple case of 1-skeletons [chung2019statistical, dakurahregistration]. Computing cycles in 1-skeletons is often done using brute force approach [tarjan1972depth]. More recently, a scalable algorithm for computing cycles in 1-skeletons and a method for computing and identifying cycles in high-order topological spaces via the Hodge Laplacian has been developed [anand2021hodge, dakurah2022modelling]. This provides an efficient algorithm for extracting cycles that retains the simplicity of graph-based approaches.
@@ 2.1 Graphs
A graph is an ordered set $\mathcal{X}=\left(V(\mathcal{X}),E(\mathcal{X}),\psi_{\mathcal{X}}\right)$, where $V(\mathcal{X})$ is the set of nodes, $E(\mathcal{X})$ is the set of edges and $\psi_{\mathcal{X}}$ is an incidence function connecting any pair of nodes of $V(\mathcal{X})$ to form edges $E(\mathcal{X})$. When the underlying graph is unambiguous from the context we usually denote the node set and vertex set as $V$ and $E$ respectively. The number of nodes and edges are also denoted as $v(\mathcal{X})$ (also denoted as p) and $e(\mathcal{X})$ respectively. Two nodes incident by the same edge are said to be adjacent, similarly two edges incident by the same node are also adjacent [bondy1991graph]. They are two important matrices associated with graphs; the adjacency matrix $\textbf{A}(\mathcal{X})$ which is a $v(\mathcal{X})\times v(\mathcal{X})$} matrix with the $a_{ij}]$ entry been the number of edges connecting node $v_{i}$ and $v_{j}$, and the incidence matrix $\textbf{B}(\mathcal{X})$ which is a $v(\mathcal{X})\times e(\mathcal{X})$ matrix with the $b_{ij}]$ entry been the number of times node $v_{i}$ is incident with edge $e_{j}$, with loops contributing two (2) [bondy1991graph, gross2005graph]. $\textbf{B}(\mathcal{X})$ is an alternate form of depicting a graph and we can derive the degree of a node by taking the appropriate column sum of $\textbf{A}(\mathcal{X})$.
@@ 2.2 Simplicial Complex and Simplicial k-chains:
A k-simplex is a k-dimensional polytope whose convex hull is made up of its k+1 nodes. A simplicial complex $\mathcal{K}$ is a set of simplices such that for any $\sigma_{1},\sigma_{2}\in\mathcal{K}$, $\sigma_{1}\cap\sigma_{2}$ is a face of both simplices; and a face of any simplex $\sigma\in\mathcal{K}$ is also a simplex in $\mathcal{K}$. For some $r_{i}\in\mathbb{R}$, the the formal sum $\sum_{i=1}^{N}r_{i}\sigma_{i}$, $\sigma_{i}\in\mathcal{K}$ is termed the simplicial k-chain. The set of simplicial k-chains with formal addition over $\mathbb{R}$ is known as an $\mathbb{R}$-module, and is used to construct boundary maps [topaz2015topological].
@@ 2.2.1 Boundary Map of Simplicial k-chains:
where $(v_{0},...,\hat{v_{i}},...,v_{k})$ is the k-1 face of $\sigma$ obtained by deleting the $v_{i}$ node. The boundary map satisfies the property that $\partial_{k}o\partial_{k+1}=0$.
@@ 2.2.2 Homology of a Simplicial Complex $\mathcal{K}$:
Two important components of the boundary map (2) are it’s kernel denoted $\mathcal{Z}_{k}=ker(\partial_{k})$, and it’s image denoted $\mathcal{B}_{k}=img(\partial_{k+1})$. These are subspaces of $\mathcal{K}_{k}$. The elements of $\mathcal{Z}_{k}$ are known as $k$-cycles and that of $\mathcal{B}_{k}$ are known as $k$-boundaries [hatcher2002algebraic]. In essence $\mathcal{Z}_{k}$ is the subspace of $\mathcal{K}_{k}$ consisting of $k$-chains that are also $k$-cycles and $\mathcal{B}_{k}$ is a subspace of $\mathcal{K}_{k}$ consisting of $k$-cycles that are also $k$-boundaries.
The set quotient $\mathcal{H}_{k}=\mathcal{Z}_{k}/\mathcal{B}_{k}$ is termed the kthkth homology-module and its an $\mathbb{R}$-module. It can be shown that $\mathcal{B}_{k}\subseteq\mathcal{Z}_{k}$ [hatcher2002algebraic, topaz2015topological]. Hence the difference, can intuitively be thought of as the failure of k-cycles in $\mathcal{K}$ to bound $(k+1)$-simplices, leading to the concept of "holes". A measure of this number of $k$-dimensional holes is termed the Betti number denoted $\beta_{k}=rank(\mathcal{H}_{k})$.
@@ 2.3 Homology of Graphs:
Graphs have natural relationship with complexes. An unoriented graph can be viewed as a 1-dimensional simplicial complex, and higher dimensional simplicial complexes are referred to as hypergraphs [lee2012persistent]. Consider the following characterization of a graph $\mathcal{X}=(V,E,w)$. Define $\mathcal{V}$ and $\mathcal{E}$ to be the node space and edge space of the graph. The formal sum over the node space
@@
Since $img(\partial_{2})=0$ [sunada2012topological], we have that the first homology module $\mathcal{H}_{1}\mathcal{(X)}=\ker(\partial_{1})$. If $\mathcal{X}$ is finite, then $\mathcal{H}_{1}\mathcal{(X)}$ has finite rank, given by the first Betti number $\beta_{1}$. Elements of $\mathcal{H}_{1}\mathcal{(X)}$ are termed 1-cycles.
@@
Let the complete graph $\mathcal{X}$ has the form $\mathcal{X}=(V,E,w)$. Since $\mathcal{X}$ is complete, it’s also connected with 1 connected component. The $0th$ Betti number $\beta_{0}$ is the number of connected components of $\mathcal{X}$. From the discussion in homology of graphs, the number of 1-cycles will be given by the $1st$ Betti number $\beta_{1}$. The Euler characteristic function of $\mathcal{X}$ is given by $\chi(\mathcal{X})=|V|-|E|=\beta_{0}-\beta_{1}$ [adler2010persistent]. The graph $\mathcal{X}$ is complete $\implies$ $|V|=p$, $|E|=\frac{|V|(|V|-1)}{2}=\frac{p(p-1)}{2}$. Hence
@@ 2.3.1 Graph Filtration
Intuitively, since $\mathcal{X}_{0}$ is complete, the number of connected components is therefore 1, as we increase threshold $\epsilon$, more edges are disconnected, increasing the number of connected components $(\beta_{0})$, and decreasing the number of cycles $(\beta_{1})$. More formally, it has been proved that in a graph $\mathcal{X}$, Betti numbers $\beta_{0}$ and $\beta_{1}$ are monotone over graph filtration on edge weights [songdechakraiwut2020topological].
@@
[adler2010persistent]. Note that $\beta_{0}(\mathcal{X}_{\epsilon})=_{0}(\mathcal{X}_{1{(w_{ij}>\epsilon)}})$ and $|V|$ is the number of nodes in $\mathcal{X}_{\epsilon}$. The function $f(|V|,\epsilon)$ is a decreasing count of the edge set over the filtration. Since the first Betti number measures the number of 1-cycles, rearranging we have:
@@ 2.3.2 Betti Curves of Graph Filtration
Given the defined weights on $\mathcal{X}$, we can induce a filtration by appealing to a graph descriptor function. A graph descriptor function function $g$ is a mapping from $\mathcal{X}$ to some $d$-dimensional space, $g:\mathcal{X}\to\mathbb{R}^{d}$. That’s, it evaluates certain attributes of $\mathcal{X}$ and embeds them in some $d$-dimensional space. For each of the subgraphs in (4), we can construct a higher dimensional path
@@
Figure 2: A schematic of a sample filtration sequence of a seven node network $\mathcal{X}$. A weighted graph (far left) is thresholded at an increasing sequence of filtration values. A graph descriptor function $g$ (count of the number of connected components) is evaluated at each filtration value. The thresholding results in a sequence of nested graphs as shown in (4). The collection of such nested sequence of graphs defines a filtration.
@@ 2.3.3 Random Graphs
Random graphs construction are mostly inspired by the Erdos-Renyi (ER) random graph [bollobas1985random]. They are two theoretical approaches to generating random graphs based of the ER model. The first approach generates a random graph with a specified set of edges, say $M$, with $v$ nodes by randomly choosing an $M$-edge graph from the subset $\binom{N}{M}$ $\frac{N!}{(N-M)!M!}$, where $N=\frac{1}{2}p(p-1)$ is the maximum number of edges in a $p$-node graph. The second approach, which is more commonly used involves assigning an edge placement probability $\pi$ to each of the $N$ possible edges. For example, assume we want to generate a random graph of order $p$ (number of nodes), we will choose each of the $N$ potential edges, run a random number generator (or equivalently flip a coin), that outputs an integer say 1 with probability $\pi$, indicating that the edge should be created, or not created otherwise. The probability space associated with this second generation process has lot of desirable properties and has been extensively studied [erdos1960evolution, gilbert1959random].
Graphs modelled after ER are often unweighted. However, complex networks will require the characterization of the heterogeneity present in edge weights. The construction of weighted random graphs has been studied as an extension of the ER model [garlaschelli2009weighted], as a generalization of the unweighted graph and also by looking at an ensemble of weighted graphs [barabasi1999emergence]. In it’s simplest form, once a random edge weight has been placed, a random weight (often modelled after a probability distribution) is assigned to the edge.